Vector A has magnitude 4 units, vector B has magnitude 6 units. The angle between A and B is 60°. Calculate the magnitude of 2A + 3B.


1
Expert's answer
2021-03-31T09:32:07-0400

"2A=2\cdot4=8"


"3B=3\cdot6=18"


"2A+3B=\sqrt{8^2+18^2+2\cdot8\cdot18\cdot\cos60°}=23.065" . Answer